# Semantic Web technologies in sensor-based personal health monitoring systems: A systematic mapping study

### v2

@@ 6.1 Data and devices
Yu et al. [110] Commercial Fitbit Withings scale Step count Body weight
@@
BG - blood glucose; BP - blood pressure; BT - body temperature; BVP - blood volume pulse; ECG - electrocardiogram; GSR - galvanic skin response; HR - heart rate; SpO2 - blood oxygen saturation; ST - skin temperature
@@ 6.2.1 Development methodologies
The use of a development methodology can streamline the process of developing Semantic Web technologies. In particular, the literature on ontology development methodologies is quite rich, with a large number of established methodologies proposed [174, 175]. There have also been several proposed approaches towards developing knowledge graphs [176] and ensuring the quality of linked data [177, 178]. Despite this, most of the systems did not report the use of a methodology in the development of the Semantic Web technologies. However, a small number of systems mentioned using a particular methodology. For example, Hadjadj and Halimi [89] used the NeOn framework [179], a scenario-based methodology for building ontologies, while Titi et al. [107] used an existing case-based ontology engineering methodology [180]. Although not a development methodology, Peral et al. [104] used the SemanTic Refinement of Ontology MAppings (STROMA) [181] approach for aligning corresponding concepts between different ontologies.
@@ 6.2.2 Development tools
Various languages, frameworks, and libraries were used to develop the systems. Among the systems that incorporated ontologies, Protégé [^33] https://protege.stanford.edu is most commonly cited as the ontology development platform of choice, mentioned in 22 of the systems. Protégé is an ontology editor that supports the latest OWL and RDF specifications. Another commonly used platform is Apache Jena [^34] https://jena.apache.org, a Java framework for building Semantic Web and Linked Data applications, mentioned in 16 of the systems. Both Protégé and Apache Jena are free and open source. When it comes to the Semantic Web languages, SWRL is the most commonly used rule language among the systems. However, Apache Jena includes a general purpose rule-based reasoner which is used by Chiang and Liang [82], Garcia-Valverde et al. [88], and Kim et al. [97]. Stavropoulos et al. [74] used SHACL to create rules, while Kordestani et al. [98] and Rhayem et al. [105] used Drools [^35] https://www.drools.org, a business rule management system. Programming languages can also be used to configure rules, as was done by Khozouie et al. [96] using Java. For queries, a majority of the systems used SPARQL, with some also using Apache Jena Fuseki, a SPARQL server, to publish their SPARQL endpoints. For storage, Mavropoulos [100] and Stavropoulos et al. [74] used GraphDB, while Spoladore et al. [106] used Stardog [^36] https://www.stardog.com. Both of these are enterprise semantic databases. Non-semantic database management systems were also used by a few of the systems, with MySQL being the most commonly cited.
@@ 6.3 Rigour of evaluation
[…]
@@
A number of systems were evaluated based on quality of service metrics. For example, Esposito et al. [86] and Vadillo et al. [108] used the Architecture-Level Modifiability Analysis (ALMA) method to evaluate the potential costs associated with modifying their systems, such as by adding more sensors. Similarly, Alti et al. [80] evaluated their system based on execution time, optimality, application’s lifetime and number of discovered services. Yu et al. [70] evaluated their system using the Chronic Care Model (CCM), an established framework for chronic care management that includes criteria such as system design, self-management support, and decision support. The systems proposed by Hristoskova et al. [92], Kim et al. [97], Mavropoulos et al. [100], Stavropoulos et al. [74], and Villarreal et al. [109] were evaluated using user studies with patients or clinicians, with Likert scales typically used to scale the user feedback. Expert validation was also used to evaluate the systems, with the aim of ensuring maximum similarity between the system output and expert opinion. This approach was taken by Ali et al. [78], El-Sappagh et al. [84], Hadjadj and Halimi [89], Hristoskova et al. [92] and Khozouie et al. [96]. Additionally, a number of systems used query-based validation, where the system is validated by checking the answers to SPARQL queries.
@@
In addition to the overall system, the system components were also evaluated. Inconsistencies in ontologies can be detected using ontological reasoners, which check whether there are contradictions in class hierarchies or class instances [4]. Reasoners such as HermiT and Pellet were used in many of the systems to evaluate the structural consistency of ontologies [79, 81, 84, 85, 86, 91, 92, 96, 107, 108, 75]. Additionally, some systems used ontology evaluation frameworks such as OntOlogy Pitfall Scanner! (OOPS!) [182], which was used by El-Sappagh et al. [84] and Zafeiropoulos et al. [75], and OQuaRE [183], which was used by Rhayem et al. [105]. Some systems also evaluated the effect of different components within the same system through ablation studies. For example, Ali et al. [77] tested the performance of their BiLSTM model for classifying healthcare data while using an ontology and without using an ontology. The results showed an increase in the accuracy of the model when combined with an ontology. Similarly, Ali et al. [78] compared the performance of their proposed ensemble deep learning model with and without feature selection. Additionally, systems that implemented ML used well-known metrics such as accuracy, precision, recall, F-score, and mean square error to evaluate the ML models.
@@ 6.4 Accessibility of research outputs
The sharing of research outputs, such as code, ontologies, knowledge graphs, and data, is a critical aspect of ensuring research is reproducible and verifiable. These resources can also be built upon by other researchers, contributing to their reuse for more efficient system development. This is severely neglected among the selected systems, with only four articles including publicly accessible links to their research outputs. Among them are Chatterjee et al. [81], who include their OWL ontology, simulated data, propositional variables, rule base, and queries as multimedia appendices. Similarly, Zafeiropoulos et al. [75] make their proposed ontology, queries, rules, code, and even research papers available via a GitHub repository. Using platforms like GitHub rather than static files has the advantage of version control, allowing researchers to manage future updates and revisions. De Brouwer et al. [83] include a link to a GitHub repository associated with the Data Analytics for Health and Connected Care ontology, which their proposed ontology extends and was developed by their research group. However, the new mBrain ontology reported in the article is not made available. In contrast, the ontology proposed by El-Sappagh et al. [84] has been published on Bioportal [^37] https://bioportal.bioontology.org, a popular repository of biomedical ontologies. However, no other research outputs, such as rules and queries, are made available.
Alti et al. [80] include a note that the data associated with their study is available by request. However, this is a suboptimal approach as it is impossible to guarantee the authors’ willingness or ability to consistently respond to such requests over time, potentially leading to prolonged delays or even a complete lack of response. Additionally, an important aspect of accessibility is ensuring that shared resources remain available in the long term. The article by Reda et al. [73] includes a link to a web portal developed as part of the system. According to the authors, the portal features a video tutorial and sample datasets for testing purposes. Unfortunately, as at the writing this article, the web portal is inaccessible and we are unable to ascertain if it was ever operational and for how long it may have been active. Finally, we note that it is understandable that researchers may be restricted in sharing participant data due to privacy concerns. A potential solution would be to seek participants’ consent in sharing their data anonymised and non-identifiable form. Table 12 indicates the links shared by researchers.
@@
6.5 Summary
This section has critically examined the quality of the selected systems, with a focus on four criteria: the data sources and devices used to collect the data; the development methodologies and tools used; the evaluation approaches and rigour; and the accessibility of research outputs. A summary list of the development tools development tools, evaluation approaches, and evaluation metrics used by the systems, can be found in Table 18 in the appendix. Though this analysis extends beyond Semantic Web technologies, we also consider several factors that are specific to Semantic Web technologies such as methodologies, languages, frameworks, and semantic databases. To summarise the section, we conduct an assessment of each system based on critical aspects related to the four quality criteria. We then discuss the results of the assessment, highlighting the aspects that are poorly addressed among the systems.
6.5.1 Quality assessment
Mirroring our assessment of how well the systems tackle the key challenges in Section 5.8, we have also evaluated the quality of the systems as reported in the corresponding research articles. We base our evaluation on the aspects of the quality criteria which are summarised in Table 13, and use the same five-point rating scale (✗, Low, Medium, High, and Very High) determined by the percentage of aspects that each system has met. The quality ratings for each system are shown in Table 19 in the appendix.
@@ 6.5.1 Quality assessment
Criteria Aspects & scoring guide
@@
4. Non-functional requirements (NFRs) have been considered and details of this have been provided (e.g. scalability, adaptability, usability, security).
@@
6.5.2 Discussion
Table 14 shows the number of systems with a particular rating for each criteria, while the combined radar chart in Figure 4 provides a visualisation of the overall quality of the systems. Separate radar charts for the individual systems are also available [^38] https://public.flourish.studio/visualisation/17845323. The accessibility of research outputs is by far the most overlooked quality criterion, with only four systems making their resources publicly accessible. With regard to the rigour of evaluation, most systems achieved either a low or medium rating. This can be primarily attributed to the fact that numerous researchers only reported the evaluation of one system component, failing to evaluate or account for the impact of other components. The potential real-world functionality of the system was another poorly addressed aspect of evaluation. Most systems were not tested with actual users; rather, case studies and scenarios were the most common approach. Additionally, most researchers overlooked the importance of evaluating their systems against external benchmarks, such as drawing comparisons with similar existing systems or seeking evaluations from domain experts.
The description of data collection methods or existing datasets was generally well done among the systems, with many of them giving adequate details regarding the number and demographics of participants or dataset records and properly citing reused datasets. However, less than half of the systems gave details of the specific devices used to collect the sensor data. A likely explanation for this is that many of the proposed systems are theoretical proposals rather than functional implementations, and therefore they were not tested on real sensor data collected from actual devices. Finally, with regard to system development, most researchers adequately reported on the tools used to develop the different components of their proposed systems. However, only a small fraction of the researchers reported the use of an existing development methodology, or else adequately described the systematic steps taken to develop each system component.
@@ 6.5.2 Discussion
11 1 2 39
Low 13 8 19 N/A
Medium 12 24 16 2
High N/A N/A 6 N/A
Very High 7 10 0 2
@@ 7 System architectures
The architecture of a system can be defined as an abstraction of the system in the form of a set of software structures needed to reason about it [184]. An important concept when discussing system architectures is the architectural style, which defines constraints on the form and structure of an architecture [185]. This is closely related to the architectural pattern, which is a reusable, well-established architectural solution to a recurring design problem [184]. As summarised in Table 4, the systems implement a range of architectural styles and patterns. This section will discuss the architectures of the systems, including how they support the achievement of the seven key challenges discussed in Section 5.

### v3

@@ 1. Polar H1 chest strap
1. Fitbit wristband 1. Step count
Yu et al. [131] Commercial 2. Withings scale 2. Body weight
@@
BG - blood glucose; BP - blood pressure; BT - body temperature; BVP - blood volume pulse; CO2 - carbon dioxide; ECG - electrocardiogram; GSR - galvanic skin response; HR - heart rate; SpO2 - blood oxygen saturation; ST - skin temperature; TVOC - total volatile organic compounds
@@ 6.2.1 Development methodologies
The use of a development methodology can streamline the process of developing Semantic Web technologies. In particular, the literature on ontology development methodologies is quite rich, with a large number of established methodologies proposed [202, 203]. There have also been several proposed approaches towards developing knowledge graphs [204] and ensuring the quality of linked data [205, 206]. Despite this, only five systems specified an existing methodology for Semantic Web technologies, with all five methodologies being ontology-focused. Hadjadj and Halimi [118] used the NeOn framework [207], a scenario-based methodology for building ontologies, while Zafeiropoulos et al. [107] used the Human-Centered Ontology engineering MEthodology (HCOME) [208]. Titi et al. [100] used an existing case-based ontology engineering methodology [209]. Similarly, Kilintzis et al. [97] followed four of the ontology construction steps recommended by Spear [210]. Although not a development methodology, Peral et al. [130] used the SemanTic Refinement of Ontology MAppings (STROMA) [211] approach for aligning corresponding concepts between different ontologies.
Additionally, two systems reported the use of existing methodologies for overall system development. Garcia-Moreno et al. [91] used the design science research methodology [212], while Vadillo et al. [101] used CommonKADS (Knowledge Acquisition and Development Systems) [213], a knowledge-based system design methodology, and its extension for multi-agent systems, MAS-CommonKADS [214]. Overall, only 16 of the systems reported or described the use of a systematic methodology, whether existing or novel, in the development of either the Semantic Web technologies, other system components, or the system as a whole.
@@ 6.2.2 Development tools
Various languages, frameworks, and libraries were used to develop the systems. When it comes to rule languages, SWRL is the most commonly used, with 24 systems either explicitly mentioning it or demonstrating SWRL syntax in code snippets [109, 111, 112, 113, 87, 114, 116, 89, 90, 118, 93, 94, 119, 124, 128, 129, 106, 83, 98, 100, 131, 107, 132, 133]. Apache Jena includes a general purpose rule-based reasoner which is used by three systems [88, 92, 122]. Other Semantic Web-based rules languages include SHACL, used by Stavropoulos et al. [99] and SPIN, used by Kilintzis et al. [97]. Although Mavropoulos et al. [126] mention using OWL to create rules, this approach is inherently limited for situation analysis and decision support because OWL lacks the expressivity for if-then rules. The authors indicate that they will implement more complex reasoning rules in future work.
Programming languages can also be used to configure rules, as is done by Khozouie et al. [121] using Java. Similarly, Fenza et al. [117] use MATLAB and Fuzzy Control Language to define their fuzzy rules, while Kordestani et al. [123] use Drools [^36] https://www.drools.org, a business rule management system. Therefore, among the 41 systems that implement rules, 25 systems use Semantic Web-based rule languages, while three use non-Semantic Web languages. Eight systems do not mention or indicate a specific formal language in which rules are defined [110, 108, 120, 95, 96, 125, 130, 104]. For queries, 26 of the systems use SPARQL, with five also using Apache Jena Fuseki, a SPARQL server, to publish their SPARQL endpoints.
Among the systems that incorporate ontologies, Protégé [^37] https://protege.stanford.edu is most commonly cited as the ontology development platform of choice, used in 27 of the systems [109, 110, 111, 112, 113, 87, 114, 88, 115, 116, 89, 90, 118, 93, 94, 120, 95, 96, 121, 122, 125, 98, 100, 101, 107, 132, 133]. Protégé is an ontology editor that supports the latest OWL and RDF specifications. Another commonly used platform is Apache Jena [^38] https://jena.apache.org, a Java framework for building Semantic Web and Linked Data applications, used in 18 of the systems [112, 113, 87, 114, 88, 115, 89, 92, 118, 94, 95, 96, 122, 83, 100, 101, 131, 132]. 27 systems report using OWL [110, 111, 112, 108, 114, 116, 89, 90, 117, 118, 94, 119, 121, 97, 122, 126, 129, 106, 83, 98, 99, 100, 101, 131, 107, 132, 133], while 16 report using RDF [108, 87, 114, 88, 115, 118, 94, 120, 97, 124, 128, 106, 100, 131, 107, 133]. Ammar et al. [108] use Solid (Social Linked Data) [^39] https://solidproject.org, a platform for developing decentralised linked data applications with the goal of enhanced privacy and data ownership. They also mention the JavaScript-based LDflex [^40] https://github.com/LDflex/LDflex and the Python-based RDFlib [^41] https://github.com/RDFLib/rdflib for linked data manipulation and querying, the latter of which is also used by Zafeiropoulos et al. [107]. These platforms and libraries are all free and open source.
For storage, Mavropoulos [126] and Stavropoulos et al. [99] use GraphDB, while Spoladore et al. [98] use Stardog [^42] https://www.stardog.com. Both of these are enterprise semantic databases. Non-semantic database management systems are also used by 11 of the systems, with SQL-based systems being more popular than NoSQL alternatives.
Four systems use MySQL [^43] https://www.mysql.com [113, 88, 100, 102], three use SQLite [^44] https://sqlite.org [116, 90, 124], and one uses PostgreSQL [^45] https://www.postgresql.org/ [125], while three systems use the NoSQL database system MongoDB [^46] https://www.mongodb.com [115, 89, 125]. A summary of all the development tools used in the selected systems is included in Table LABEL:development-evaluation-summary in the appendix.
@@ 6.3 Rigour of evaluation
A variety of evaluation approaches are used by the selected systems. The most common approach is case-based evaluation. 21 systems were evaluated through use case scenarios, which generally describe the sequence of events when a user interacts with the system [108, 87, 88, 115, 116, 89, 92, 118, 96, 121, 123, 127, 128, 98, 99, 101, 104, 107, 132, 133, 103]. Nine systems were evaluated using case studies, which are similar to use case scenarios but are more extensive and detailed [109, 113, 90, 117, 93, 125, 129, 130, 102, 105]. 12 systems were evaluated by running user studies with real users [112, 115, 90, 91, 119, 120, 97, 122, 126, 99, 102, 104]. Among these systems, three used Likert scales to measure user feedback [122, 126, 99]. Additionally, 16 compared their systems with existing ones, showing how their results performed against the state of the art using a set criteria [111, 112, 116, 120, 95, 126, 83, 105].
@@
21 systems were evaluated based on non-functional requirements. For example, Esposito et al. [90] and Vadillo et al. [101] used the Architecture-Level Modifiability Analysis (ALMA) method to evaluate the potential costs associated with modifying their systems, such as by adding more sensors. Similarly, Alti et al. [113] evaluated their system based on execution time, optimality, application’s lifetime and number of discovered services. Domain-specific quality metrics were also considered. Yu et al. [104] evaluated their system using the Chronic Care Model (CCM), an established framework for chronic care management that includes criteria such as system design, self-management support, and decision support.
Additionally, five systems used simulation as a means to investigate the system functionality. For example, Akhtar et al. [109] used Netlogo, a multi-agent modelling platform, to simulate the use of their system. Chiang and Liang [88] used a fuzzy logic simulation tool to validate their fuzzy inference module. Ivașcu and Negru [95] simulated the system functionality by using each subject in the dataset as the target user, while Martella et al. [125] used a testbed environment to simulate user scenarios and test how well the system responds to varying workloads. Finally, Reda et al. [106] used a web portal with sample data for testing purposes. Expert validation was also used to evaluate the systems, with the aim of ensuring maximum similarity between the system output and expert opinion. This approach was taken by Ali et al. [111], El-Sappagh et al. [116], Hadjadj and Halimi [118], Hristoskova et al. [119] and Khozouie et al. [121]. Additionally, 10 systems used query-based validation, where the system is validated by checking the answers to SPARQL queries [112, 114, 115, 116, 118, 97, 122, 100, 131, 107].
@@
In addition to the overall system, the system components were also evaluated. Inconsistencies in ontologies can be detected using ontology reasoners, which check whether there are contradictions in class hierarchies or class instances [4]. Reasoners were used in 11 of the systems to evaluate the structural consistency of ontologies. Eight systems used Pellet [112, 89, 90, 94, 119, 121, 100, 101, 107], one system used HermiT [114], and one system reported using both Pellet and HermiT [116]. Additionally, three systems used ontology evaluation frameworks, namely OntOlogy Pitfall Scanner! (OOPS!) [215], which was used by El-Sappagh et al. [116] and Zafeiropoulos et al. [107], and OQuaRE [216], which was used by Rhayem et al. [83]. Two systems also evaluated the effect of different components within the same system through ablation studies. Ali et al. [110] tested the performance of their BiLSTM model for classifying healthcare data while using an ontology and without using an ontology. The results showed an increase in the accuracy of the model when combined with an ontology. Similarly, Ali et al. [111] compared the performance of their proposed ensemble deep learning model with and without feature selection. Finally, 13 systems, mainly those that implement ML, used well-known metrics such as accuracy, precision, recall, F-score, and mean square error [110, 111, 112, 113, 91, 92, 120, 95, 122, 126, 83, 104, 107, 132].
@@ 6.4 Accessibility of research outputs
The sharing of research outputs, such as code, ontologies, knowledge graphs, and data, is a critical aspect of ensuring research is reproducible and verifiable. These resources can also be built upon by other researchers, contributing to their reuse for more efficient system development. This is severely neglected among the selected systems, with only six articles including links to their research outputs. Among them are Chatterjee et al. [114], who include their OWL ontology, simulated data, propositional variables, rule base, and queries as multimedia appendices. Using platforms like GitHub rather than static files has the advantage of version control, allowing researchers to manage future updates and revisions. Three systems take this approach. The system proposed by Bampi et al. [87] is available on GitHub, including the backend, frontend, ontology, rules, and queries. Similarly, Zafeiropoulos et al. [107] make their proposed ontology, queries, rules, code, and even research papers available via a GitHub repository. De Brouwer et al. [115] include a link to a GitHub repository associated with the Data Analytics for Health and Connected Care ontology, which their proposed ontology extends and was developed by their research group. However, the new mBrain ontology reported in the article is not made available. In contrast, the ontology proposed by El-Sappagh et al. [116] has been published on Bioportal [^47] https://bioportal.bioontology.org, a popular repository of biomedical ontologies. However, no other research outputs, such as rules and queries, are made available.
An important consideration when sharing system resources is ensuring their long-term accessibility. For instance, although Kilintzis et al. [97] and Reda et al. [106] provide links to their systems, the web pages were unavailable at the time of writing this article. Kilintizis et al. [97] stated that the link made their ontology available either for reuse or review purposes, while the link from Reda et al. [106] was to a portal with a video tutorial and sample datasets for testing purposes. As these resources are now inaccessible, we are unable to ascertain if they were ever operational and for how long they may have been active. Additionally, while GitHub repositories allow for easy accessibility and future updates, they can also be deleted or made private. A good alternative is Zenodo [^48] https://zenodo.org, a general-purpose research repository designed for long-term preservation by ensuring that published records can be updated but not unpublished. Further, Zenodo provides persistent digital object identifiers (DOIs) for each upload, ensuring easy referencing.
Alti et al. [113] and Garcia-Moreno et al. [91] include a note that the data associated with their studies are available upon request. However, this is a suboptimal approach as it is impossible to guarantee the authors’ willingness or ability to consistently respond to such requests over time, potentially leading to prolonged delays or even a complete lack of response. Publishers can mitigate this through well-defined data availability policies. Finally, we note that researchers may be restricted from sharing participant data due to privacy concerns. A potential solution would be to seek participants’ consent in sharing their data anonymised and non-identifiable form. Table 13 indicates the links shared by researchers.
@@
Bampi et al. [87] GitHub repository https://github.com/mbampi/aal-system
@@
Kilintzis et al. [97] Web page - currently inaccessible http://lomi.med.auth.gr/ontologies/WELCOME_entities
@@
6.5 Quality assessment
Mirroring our assessment of how well the systems tackle the key challenges in Section 5.9, we have also evaluated the quality of the systems as reported in the corresponding research articles. We base our evaluation on the aspects of the quality criteria which are summarised in Table LABEL:quality-critera, and use the same four-point rating scale (✗, Low, Medium, and High) determined by the percentage of aspects that each system has met. The quality ratings for each system are shown in Table LABEL:quality_summary in the appendix.
@@ 6.5 Quality assessment
Criteria Aspects and scoring guide
@@
4. Non-functional requirements have been considered and details of this have been provided (e.g. scalability, adaptability, usability, security).
@@
Table 15 shows the number of systems with a particular rating for each criteria, while the combined radar chart in Figure 4 provides a visualisation of the overall quality of the systems. Separate radar charts for the individual systems are also available [^49] https://public.flourish.studio/visualisation/17845323.
@@
11 1 2 42
Low 14 10 20 N/A
Medium 13 25 20 3
High 10 12 6 3
@@
6.6 Summary
This section has critically examined the quality of the selected systems, with a focus on four criteria: the data sources and devices used to collect the data; the development methodologies and tools used; the evaluation approaches and rigour; and the accessibility of research outputs. A summary list of the development tools development tools, evaluation approaches, and evaluation metrics used by the systems, can be found in Table LABEL:development-evaluation-summary in the appendix. Though this analysis extends beyond Semantic Web technologies, we also consider several factors that are specific to Semantic Web technologies such as methodologies, languages, frameworks, and semantic databases. To summarise the section, we discuss the results of the quality assessment, highlighting the aspects that are poorly addressed among the systems.
The accessibility of research outputs is by far the most overlooked quality criterion, with only six systems making their resources publicly accessible. With regard to the rigour of evaluation, 20 systems achieved a low rating, with another 20 achieving a medium rating. This can be primarily attributed to the fact that numerous researchers only reported the evaluation of one system component, failing to evaluate or account for the impact of other components. The potential real-world functionality of the system was another poorly addressed aspect of evaluation. Most systems were not tested with actual users, with only 12 systems being user validated. Case studies and scenarios were the most common approach, used in 29 of the systems. Additionally, most researchers (32 systems) overlooked the importance of evaluating their systems against external benchmarks, such as drawing comparisons with similar existing systems or seeking evaluations from domain experts.
The description of data collection methods or existing datasets was generally well done among the systems, with 30 of them giving adequate details regarding the number and demographics of participants or dataset records and properly citing reused datasets. However, only 19 of the systems gave details of the specific devices used to collect the sensor data. A likely explanation for this is that many of the proposed systems are conceptual proposals rather than functional implementations, and therefore they were not tested on real sensor data collected from actual devices. Finally, with regard to system development, 36 papers comprehensively reported on the tools used to develop the different components of their proposed systems. However, only 16 of the papers reported the use of an existing development methodology, or else adequately described the systematic steps taken to develop each system component.
@@ 7 System architectures
The architecture of a system can be defined as an abstraction of the system in the form of a set of software structures needed to reason about it [217]. An important concept when discussing system architectures is the architectural style, which defines constraints on the form and structure of an architecture [218]. This is closely related to the architectural pattern, which is a reusable, well-established architectural solution to a recurring design problem [217]. As summarised in Table LABEL:selected-systems, the systems implement a range of architectural styles and patterns. This section will discuss the architectures of the systems, including how they support the achievement of the seven key challenges discussed in Section 5.